# How to discretize continuous state-action spaces in Q-learning: A symbolic control approach

### v2

@@ VII Controller refinement and Q-learning for symbolic model
VII-A Controller refinement
Q-learning is a reinforcement learning algorithm that is known for its off-policy nature. This key characteristic provides the designer with the flexibility to evaluate and compare both the resultant maximally and minimally optimal policies in an offline manner. By evaluating and analysing the performance of these policies, the designer can choose the one that best fits the requirements of the control problem at hand.
VII-B Q-learning for symbolic model
The proposed Q-learning algorithm is shown in Algorithm 3:
@@ VII-B Q-learning for symbolic model
1:  Input: Environment with continuous state space $\mathcal{S}$, continuous action space $\mathcal{A}$. Define the Lipshitz constant $L_{f}$ and $L_{g}$, the learning rate $\alpha$ and the discount factor $\gamma$. Choose an arbitrary Q-value precision $\varepsilon$. Initialize $Q_{min}$ and $Q_{max}$ satisfying Assumption 4.
2:  Output: Learned $Q_{min}$ and $Q_{max}$-values for all symbolic states and action samples.
@@
11:        $Q_{min}(s,a)\leftarrow Q_{min}(s,a)+\alpha(\underline{g}+\gamma\min\limits_{s^{\prime}\in\Delta(s,a)}\max\limits_{a^{\prime}\in\mathcal{A}_{q}(s^{\prime})}Q_{min}(s^{\prime},a^{\prime}))$
12:        $Q_{max}(s,a)\leftarrow Q_{max}(s,a)+\alpha(\underline{g}+\gamma\max\limits_{s^{\prime}\in\Delta(s,a)}\max\limits_{a^{\prime}\in\mathcal{A}_{q}(s^{\prime})}Q_{max}(s^{\prime},a^{\prime}))$
@@
Note that in the present paper, the reward function is redefined for each discrete state action pair. This permits to apply Q-learning in a fully discrete setting, without any involvement of the model with continuous state action spaces.

### v3

@@ VII Controller refinement and Q-learning for symbolic model
Q-learning is an off-policy reinforcement learning algorithm, offering the advantage of evaluating and comparing maximally and minimally optimal policies in an offline manner. This flexibility allows the designer to select the policy that best suits the control problem’s requirements.
The proposed Q-learning algorithm is shown in Algorithm 3. Once the learning is established, any of the Q-tables $\underline{q}$ and $\overline{q}$ can be used to refine the learned strategy into a controller for the original system $\Sigma$ in (1). Formally, a controller for the original discrete-time system $\Sigma$ in (1) is a map $\pi:\mathcal{S}\rightarrow\mathcal{A}$ satisfying $\pi(\xi)\in\mathcal{A}(\xi)$ for all $\xi\in\mathcal{S}$. Now given the results of the $Q$ tables $\underline{q}$ and $\overline{q}$, two control strategies are generated:
  • A lower control strategy $\underline{\pi}:\mathcal{S}\rightarrow\mathcal{A}$ defined formally as follows: for $\xi\in\mathcal{S}$ such that $\xi\in s$, $s\in\mathcal{S}_{\mathrm{D}}$, we have $\underline{\pi}(\xi)\in\arg\max\limits_{a}\underline{q}(s,a)$;
  • An upper control strategy $\overline{\pi}:\mathcal{S}\rightarrow\mathcal{A}$ defined formally as follows: for $\xi\in\mathcal{S}$ such that $\xi\in s$, $s\in\mathcal{S}_{D}$, we have $\overline{\pi}(\xi)\in\arg\max\limits_{a}\overline{q}(s,a)$;
One may be in the case where there is a disagreement between the policies generated by $\underline{q}$ and $\overline{q}$, if they disagree, one can refine either using $\underline{q}$ or $\overline{q}$ while preserving $\varepsilon$-approximate optimality, ensuring (39). Remark 5 in the Appendix 2 highlight other differences between algorithm 3 and the algorithm in [13].
@@
1:  Input: Environment with continuous state space $\mathcal{S}$, continuous action space $\mathcal{A}$. Define the Lipshitz constant $L_{f}$ and $L_{g}$, the learning rate $\alpha$ and the discount factor $\gamma$. Choose an arbitrary Q-value precision $\varepsilon$. Initialize $\underline{q}$ and $\overline{q}$ satisfying Assumption 4.
2:  Output: Learned $\underline{q}$ and $\overline{q}$-values for all symbolic states and action samples.
@@
11:        $\underline{q}(s,a)\leftarrow\underline{q}(s,a)+\alpha(\underline{g}+\gamma\min\limits_{s^{\prime}\in\Delta(s,a)}\max\limits_{a^{\prime}\in\mathcal{A}_{q}(s^{\prime})}\underline{q}(s^{\prime},a^{\prime}))$
12:        $\overline{q}(s,a)\leftarrow\overline{q}(s,a)+\alpha(\underline{g}+\gamma\max\limits_{s^{\prime}\in\Delta(s,a)}\max\limits_{a^{\prime}\in\mathcal{A}_{q}(s^{\prime})}\overline{q}(s^{\prime},a^{\prime}))$